A population consists of the four numbers 3,7,11,15,the variance of the population is


1
Expert's answer
2021-01-21T02:40:30-0500

We know the formula for variance of population such as , "\sigma ^2=\frac{\sum_{i=1}^{N} (x_i-\mu)^2}{N}"

where "x_i=" value of "i" th element

"\mu =" population mean

"N=" population size

Here the given population data are "3,7,11,15"

Now "\mu =\frac{(3+7+11+15)}{4}=9" , "N=4"

and "\sum_{i=1}^{N}(x_i-\mu)^2=(3-9)^2+(7-9)^2+(11-9)^2+(15-9)^2=80"

"\therefore \sigma^2=\frac {80}{4}=20"

Hence required variance of the population is "=20"